The length of the top of a table is 5 m greater than the width/ The area is 104m^2. What is the dimension.
Answer by Boreal(15235)

width=x
length=x+5
x*(x+5)=104
x^2+5x-104=0
(x+13)(x-8)=0
x= -13, 8;;only 8 counts
width 8 m
length 13 m